# Understanding User Mental Models in AI-Driven Code Completion Tools: Insights from an Elicitation Study

### v2

@@ 4.4 Data collection
The researchers analyzed the transcripts systematically using an inductive thematic analysis (Braun and Clarke, 2006). According to the thematic analysis protocol, the evaluators performed the following 6 steps: familiarizing with the data, generating codes, mapping codes that share similar meanings, reviewing potential themes, defining and naming themes, and producing a report. The eight guiding questions structured the analysis, with the final themes answering each question.
@@ 5.1 UI for code completion
  • Hybrid (6 groups): a basic behavior could involve a proactive AI strategy that intervenes only when there is a well-defined context (e.g., a long code already written) and when the suggestion is almost taken for granted and difficult to discard by the user (e.g., completion of a control structure). On the other hand, manual intervention would compensate the user for the lack of automatic AI intervention, provided that the user is sure that they require AI support (“Proactive when it’s really obvious and useful, or when the user explicitly requests it” – G8). Furthermore, several participants specify that, in general, suggestions should be generated only if there is enough coding context available (4 groups) (“It’s better not to try immediately, but take from context” – G2). This was also consistent with the perceived necessity of very precise, context-sensitive recommendations (Asaduzzaman et al., 2014).
@@ 6.2 Visualization preferences (RQ1)
The visualization mechanism of suggestions impacts the usability and acceptance of CCTs. The participants proposed different visualization strategies, each with pros and cons. As shown in Section˜5.1 (‘How to show suggestions?’) and summarized in Fig.˜2, the majority of groups (5/8) preferred inline suggestions, as well as suggestions displayed in sidebars (7/8). Having inline suggestions can be useful to avoid losing the coding focus and allow the user to accept or reject the suggestion immediately. However, the inline solution is well-suited for short and concise suggestions that do not interfere too much with the code manually written by the users; on the contrary, in the case of long and complex suggestions, it may be better to have separate sidebars to limit visual clutter. A further level of control would be having users manually invoke popups containing the suggestion by clicking an icon shown close to the handwritten code (proposed by 2 groups). Moreover, a chatbot in a separate window could be used as an alternative way to suggest code on demand (proposed by 2 groups), with all the attached benefits of using conversational agents for code generation (Ross et al., 2023).
@@ 6.4 Preferences for explanation delivery (RQ2)
Participants generally considered explanations useful, but the answers to the question of what should be contained in an explanation revealed a need for special care in their design. The concept of what defines a good explanation is not trivial and is an aspect that is widely studied in the literature (Miller, 2019; Holzinger et al., 2020). Moreover, even well-designed explanations should be evaluated with users to measure their quality (Holzinger et al., 2020). This elicitation study can be a precious source of information in understanding the desiderata of explanations in CCTs from the developers’ point of view. A key finding is that explanations should be classified into two categories: why-explanations and what-explanations. Why-explanations address the rationale or purpose behind the generated code, whereas what-explanations concentrate on the functionality and technical intricacies of the code suggestion. The content should, however, be minimal, with the possibility to obtain more details when requested. This might translate into the CCT by initially showing a brief why-explanation, providing a bird-eye view of what the generated code is for. A more thorough, on-demand explanation would instead detail the code nuances and technical aspects to help the developer get a deeper understanding of the suggestion, possibly even with flowcharts. Detailed explanations should also include references to trusted sources, such as documentation snippets, book pieces, etc., and even examples of use to enhance understanding.
Regarding where to show explanations, participants typically preferred separate elements (e.g., popups, sidebars), mainly because they did not want the explanations to be invasive and occupy space within the coding area. However, inline explanations may be accepted as comments close to the suggested code as long as their length is very limited. However, separate windows may allow developers to initiate an explanation process with a conversational agent, similar to what is possible to achieve with tools such as ChatGPT (OpenAI, 2024a).
@@ 9 Conclusion
Limitations of this study mostly regard the convenience sampling methodology that was adopted, which led to a sample of very young participants, with limited to no expertise in companies. Therefore, despite the large sample size, the findings of this study are not directly generalizable to professional developers. To address this limitation we plan to conduct usablity studies with different categories of users. Such studies will be initially conducted in a laboratory setting, and then in-the-wild within an IT company; this will allow us to measure the effectiveness and usability of ATHENA compared to other popular CCTs. Moreover, we will observe how different features of a CCT can affect the usability and performance of users with varying degrees of expertise with different categories of users. Future work will also entail further developing our CCT and extending its functionalities by implementing the remaining features that emerged from the elicitation study. Moreover, we will explore the use of different LLMs as the AI code completion engine of the tool; in fact, to generate code suggestions and explanations ATHENA employs OpenAI’s GPT-4o mini, an LLM which has a limited size but performs well on coding problems (OpenAI, 2024b). However, different models may prove more suitable for use in corporate settings. For instance, a non-commercial solution such as Meta’s open-source LLM LLama (Meta, 2024) could be deployed on a company’s server to prevent sharing confidential data, such as project repositories, with third parties. Furthermore, models of varying sizes could be contemplated in light of a tradeoff between costs and performance.

### v3

@@ 4.4 Data collection
The researchers analyzed the transcripts systematically using an inductive thematic analysis (Braun2006Using). According to the thematic analysis protocol, the evaluators performed the following 6 steps: familiarizing with the data, generating codes, mapping codes that share similar meanings, reviewing potential themes, defining and naming themes, and producing a report. The eight guiding questions structured the analysis, with the final themes answering each question.
@@ 5.1 UI for code completion
  • Hybrid (6 groups): a basic behavior could involve a proactive AI strategy that intervenes only when there is a well-defined context (e.g., a long code already written) and when the suggestion is almost taken for granted and difficult to discard by the user (e.g., completion of a control structure). On the other hand, manual intervention would compensate the user for the lack of automatic AI intervention, provided that the user is sure that they require AI support (“Proactive when it’s really obvious and useful, or when the user explicitly requests it” – G8). Furthermore, several participants specify that, in general, suggestions should be generated only if there is enough coding context available (4 groups) (“It’s better not to try immediately, but take from context” – G2). This was also consistent with the perceived necessity of very precise, context-sensitive recommendations (asaduzzaman2014cscc).
@@ 6.2 Visualization preferences (RQ1)
The visualization mechanism of suggestions impacts the usability and acceptance of CCTs. The participants proposed different visualization strategies, each with pros and cons. As shown in Section˜5.1 (‘How to show suggestions?’) and summarized in Fig.˜2, the majority of groups (5/8) preferred inline suggestions, as well as suggestions displayed in sidebars (7/8). Having inline suggestions can be useful to avoid losing the coding focus and allow the user to accept or reject the suggestion immediately. However, the inline solution is well-suited for short and concise suggestions that do not interfere too much with the code manually written by the users; on the contrary, in the case of long and complex suggestions, it may be better to have separate sidebars to limit visual clutter. A further level of control would be having users manually invoke popups containing the suggestion by clicking an icon shown close to the handwritten code (proposed by 2 groups). Moreover, a chatbot in a separate window could be used as an alternative way to suggest code on demand (proposed by 2 groups), with all the attached benefits of using conversational agents for code generation (Ross2023Programmers).
@@ 6.4 Preferences for explanation delivery (RQ2)
Participants generally considered explanations useful, but the answers to the question of what should be contained in an explanation revealed a need for special care in their design. The concept of what defines a good explanation is not trivial and is an aspect that is widely studied in the literature (Miller2019Explanation; Holzinger2020Measuring). Moreover, even well-designed explanations should be evaluated with users to measure their quality (Holzinger2020Measuring). This elicitation study can be a precious source of information in understanding the desiderata of explanations in CCTs from the developers’ point of view. A key finding is that explanations should be classified into two categories: why-explanations and what-explanations. Why-explanations address the rationale or purpose behind the generated code, whereas what-explanations concentrate on the functionality and technical intricacies of the code suggestion. The content should, however, be minimal, with the possibility to obtain more details when requested. This might translate into the CCT by initially showing a brief why-explanation, providing a bird-eye view of what the generated code is for. A more thorough, on-demand explanation would instead detail the code nuances and technical aspects to help the developer get a deeper understanding of the suggestion, possibly even with flowcharts. Detailed explanations should also include references to trusted sources, such as documentation snippets, book pieces, etc., and even examples of use to enhance understanding.
Regarding where to show explanations, participants typically preferred separate elements (e.g., popups, sidebars), mainly because they did not want the explanations to be invasive and occupy space within the coding area. However, inline explanations may be accepted as comments close to the suggested code as long as their length is very limited. However, separate windows may allow developers to initiate an explanation process with a conversational agent, similar to what is possible to achieve with tools such as ChatGPT (ChatGPT).
@@ 9 Conclusion
Limitations of this study mostly regard the convenience sampling methodology that was adopted, which led to a sample of very young participants, with limited to no expertise in companies. Therefore, despite the large sample size, the findings of this study are not directly generalizable to professional developers. To address this limitation we plan to conduct usablity studies with different categories of users. Such studies will be initially conducted in a laboratory setting, and then in-the-wild within an IT company; this will allow us to measure the effectiveness and usability of ATHENA compared to other popular CCTs. Moreover, we will observe how different features of a CCT can affect the usability and performance of users with varying degrees of expertise with different categories of users. Future work will also entail further developing our CCT and extending its functionalities by implementing the remaining features that emerged from the elicitation study. Moreover, we will explore the use of different LLMs as the AI code completion engine of the tool; in fact, to generate code suggestions and explanations ATHENA employs OpenAI’s GPT-4o mini, an LLM which has a limited size but performs well on coding problems (2024GPT4oMini). However, different models may prove more suitable for use in corporate settings. For instance, a non-commercial solution such as Meta’s open-source LLM LLama (2024LLama3-2) could be deployed on a company’s server to prevent sharing confidential data, such as project repositories, with third parties. Furthermore, models of varying sizes could be contemplated in light of a tradeoff between costs and performance.